A 70.0 g piece of metal at 80.0oC is placed in 1.0 x 102 g of water at 22.0oC contained in a calorimeter. The metal and water come to the same temperature at 24.6oC. How much heat (in kJ) did the metal give up to the water?


1
Expert's answer
2021-04-06T04:35:19-0400

Cp(H2O) = 4,18 "\frac{J}{g\times K}"

"\Delta"T(H2O) = (24.6 - 22) = 2.6 oC = 2.6 oK

Q = 4.18"\times"2.6"\times"100 = 1086.8 J = 1.09 kJ

Answer: Q = 1.09 kJ.